# Holographic Intelligence Surface Assisted Integrated Sensing and Communication

### v1

@@ VI Conclusions
This paper introduced a novel concept to ISAC systems by proposing the HISAC system. Through the continuous-aperture array fabrication of both transmitter and receiver, the HISAC system presented unique challenges in design. To address these challenges, a continuous-discrete transformation technique based on the Fourier transform was employed, facilitating the conversion of continuous pattern design into discrete beamforming design. Subsequently, a joint transmit-receive beamforming optimization problem was formulated to balance multi-target sensing performance with multi-user communication requirements. An AO-based algorithm was developed to efficiently tackle the non-convex optimization problem with coupled variables. Numerical results demonstrated superior performance compared to traditional discrete-array-based ISAC systems, achieving significantly enhanced sensing capabilities while maintaining predetermined communication performance levels.
-A Proof of Proposition 1
The expression of the scalar Green’s function on the far-field is given by
$\centering G(\textbf{r},\textbf{s})=\frac{e^{j\kappa\|\textbf{r}-\textbf{s}\|}}{4\pi\|\textbf{r}-\textbf{s}\|},$ (-A.1)
[…]
$=\sqrt{\left(r\sin{\theta}\cos{\psi}-s_{x}\right)^{2}+\left(r\sin{\theta}\sin{\psi}-s_{y}\right)^{2}+r^{2}\cos{\theta}}$
$\approx r-\sin{\theta}\left(s_{x}\cos{\psi}+s_{y}\sin{\psi}\right).$
By substituting (-A.2) into (-A.1), the Green’s function on the far field is approximated as
$\centering G_{p^{\prime}}=\frac{e^{j\kappa r}}{4\pi r}e^{-j\kappa s_{x}\sin{\theta}\cos{\psi}}e^{-j\kappa s_{y}\sin{\theta}\sin{\psi}}.$ (-A.3)
Combining (-A.3) and (9), the Green’s function Fourier transform $f_{2D}^{n}\left(G_{p^{\prime}}\right)$, is given by
$\centering f_{2D}^{n}\left(G_{p^{\prime}}\right)=\int_{D}\frac{e^{j\kappa r}}{4\pi r}e^{-j\kappa s_{x}\sin{\theta}\cos{\psi}}e^{-j\kappa s_{y}\sin{\theta}\sin{\psi}}\Psi_{n}(p)\mathrm{d}p,$ (-A.4)
$\centering\Psi_{n}(p)=\frac{1}{\sqrt{A_{T}}}e^{-j2\pi\left(\frac{n_{x}}{L_{x}}\left(s_{x}-\frac{L_{x}}{2}\right)\right)}e^{-j2\pi\left(\frac{n_{y}}{L_{y}}\left(s_{y}-\frac{L_{y}}{2}\right)\right)}.$ (-A.5)
Let $\kappa_{x}^{n}=\kappa\left(\sin{\theta}\cos{\psi}+\lambda\frac{n_{x}}{L_{x}}\right)$ and $\kappa_{y}^{n}=\kappa\left(\sin{\theta}\sin{\psi}+\lambda\frac{n_{y}}{L_{y}}\right)$ denote the wavenumbers along the $x$-axis and $y$-axis, respectively. Combining (-A.4) with (-A.5), the Green’s function Fourier transform for arbitrary position $(r,\theta,\psi)$ is calculated as follows:
$\centering f_{2D}^{n}\left(G_{p^{\prime}}\right)=$ $\frac{e^{j\kappa r}}{4\pi r\sqrt{A_{T}}}e^{j(n_{x}+n_{y})\pi}$ (-A.6)
$\int_{L_{x}}e^{-j\kappa_{x}s_{x}}\mathrm{d}s_{x}\int_{L_{y}}e^{-j\kappa_{y}s_{y}}\mathrm{d}s_{y}$
$=$ $\frac{e^{j\kappa r}}{4\pi r\sqrt{A_{T}}}e^{j(n_{x}+n_{y})\pi}$
$\left[\frac{e^{-j\kappa_{x}^{n}s_{x}}}{-j\kappa_{x}^{n}}\right]_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}\left[\frac{e^{-j\kappa_{y}^{n}s_{y}}}{-j\kappa_{y}^{n}}\right]_{-\frac{L_{y}}{2}}^{\frac{L_{y}}{2}}$
$=$ $\frac{e^{j\kappa r}}{4\pi r\sqrt{A_{T}}}e^{j(n_{x}+n_{y})\pi}$
$\frac{L_{x}\sin{\left(\kappa_{x}^{n}\frac{L_{x}}{2}\right)}}{k_{x}^{n}\frac{L_{x}}{2}}\frac{L_{y}\sin{\left(\kappa_{y}^{n}\frac{L_{y}}{2}\right)}}{k_{y}^{n}\frac{L_{y}}{2}}$
Finally, we get an expression of the Green’s function Fourier transform for arbitrary position $(r,\theta,\psi)$:
$\centering f_{2D}^{n}\left(G_{p^{\prime}}\right)=$ $\frac{e^{j\kappa r}\sqrt{A_{T}}}{4\pi r}e^{j(n_{x}+n_{y})\pi}$ (-A.7)
$\mathrm{sinc}{\left(\kappa_{x}^{n}\frac{L_{x}}{2}\right)}\mathrm{sinc}{\left(\kappa_{y}^{n}\frac{L_{y}}{2}\right)}.$
@@
-B Proof of Proposition 2
$\centering\boldsymbol{\hat{w}}_{k}=\left(\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k}\right)^{-\frac{1}{2}}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k},~{}k=1,...,K,$ (-B.1)
and the beamforming matrix $\boldsymbol{\hat{W}}_{r}$ for sensing is choosing by
$\centering\boldsymbol{\hat{W}}_{r}\boldsymbol{\hat{W}}_{r}^{H}=\boldsymbol{\hat{R}}-\sum_{k=1}^{K}\boldsymbol{\hat{R}}_{k}.$ (-B.2)
To show the beamforming solution $\boldsymbol{\hat{W}}_{c}$ and $\boldsymbol{\hat{W}}_{r}$ obtained by using (-B.1) and (-B.2) are solutions of (23), we need to prove that the reconstructed matrices $\boldsymbol{\bar{R}}=\boldsymbol{\hat{W}}_{c}\boldsymbol{\hat{W}}_{c}^{H}+\boldsymbol{\hat{W}}_{r}\boldsymbol{\hat{W}}_{r}^{H}$, $\boldsymbol{\bar{R}}_{k}=\boldsymbol{\hat{w}}_{k}\boldsymbol{\hat{w}}_{k}^{H}$, for $k=1,...,K$, remain the same optimal and satisfy all the constraints of problem (23).
First, since the objective function in (23) is only determined by the optimal solution $\boldsymbol{\hat{R}}$, the objective function in (23) satisfies. Obviously, the reconstructed matrices $\boldsymbol{\bar{R}}_{1},...,\boldsymbol{\bar{R}}_{K}$ are rank-one, satisfying (23g).
Then, we only need to prove that matrices $\boldsymbol{\bar{R}}_{1},...,\boldsymbol{\bar{R}}_{K}$, satisfy constraints (23b), (23c). According to (-B.1), we have the following equivalent formula,
$\centering\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{w}}_{k}\boldsymbol{\hat{w}}_{k}^{H}\boldsymbol{f}_{k}$ $=\boldsymbol{f}_{k}^{H}\left(\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k}\right)^{-1}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k}\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k}$ (-B.3)
$\equiv\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k},~{}k=1,...,K,$
which indicates the SINR constraints (23b) for user communication are satisfied. Since the reconstructed solutions $\boldsymbol{\bar{R}}_{k}$, for $k=1,...,K$, satisfy $\boldsymbol{f}_{k}^{H}\boldsymbol{\bar{R}}_{k}\boldsymbol{f}_{k}\equiv\boldsymbol{f}_{k}^{H}\boldsymbol{\hat{R}}_{k}\boldsymbol{f}_{k}\geq 0$, we have $\boldsymbol{\bar{R}}_{k}\succeq 0,~{}\forall k$, satisfying (23c).
This concludes the proof.

### v2

@@ VI Conclusion and Discussion
This paper introduced a novel concept to ISAC systems by proposing the HISAC system. Through the continuous-aperture array fabrication of both transmitter and receiver, the HISAC system presented unique challenges in design. To address these, we developed a Fourier-based continuous–discrete transformation that maps the continuous pattern into a finite-dimensional wavenumber domain, enabling compatibility with traditional discrete-array ISAC frameworks Based on this reformulation, a joint transmit-receive beamforming problem was formulated to balance multi-target sensing performance with multi-user communication requirements. An AO-based algorithm was developed to efficiently tackle the non-convex optimization with coupled variables. Numerical results demonstrated superior performance compared to traditional discrete-array-based ISAC systems, achieving significantly enhanced sensing capabilities while maintaining predetermined communication performance levels.
VI-A Discussion
VI-A1 Hardware Implementation Challenges in HIS Systems
This work focuses on the algorithmic and theoretical modeling of continuous-aperture HIS. However, practical implementation introduces several hardware-level challenges that must be addressed before real-world deployment. The proposed HISAC framework assumes a fully-digital HIS architecture, in which each beam, corresponding to a Fourier-domain coefficient, is driven by a dedicated RF chain. However, implementing such a system in practice entails significant hardware complexity, particularly due to the delicate RF chains. To bridge the gap between theoretical models and practical systems, future implementations may consider hybrid HIS architectures, which offer a trade-off between flexibility and complexity. In addition, hardware imperfections such as phase quantization, nonlinearities, and finite control resolution should be carefully modeled to understand their impact on system performance. These non-idealities primarily distort the effective HIS beampattern and can be incorporated into the proposed framework through uncertainty-aware constraints or robust optimization formulations.
VI-A2 Scalability in Multi-HIS and Large-Scale Networks
The proposed HISAC system already adopts a dual-HIS configuration, in which separate transmit and receive HISs are deployed to support full-duplex ISAC functionality. However, extending the framework to cell-free MIMO or multistatic sensing, which requires the cooperation among multiple base stations/HISs, has yet to be explored. In such distributed settings, inter-HIS synchronization, mutual interference suppression, and coordinated beam scheduling become essential. As the number of cooperative HIS nodes increases, the dimensionality of control variables grows accordingly, which substantially increases the computational complexity of joint optimization. For large-scale deployments involving multiple HISs, future research may investigate low-rank structural modeling, learning-assisted beam coordination strategies, and distributed optimization algorithms. These approaches can facilitate scalable and adaptive system operation while maintaining global performance constraints.
VI-A3 System-Level Design Based on the HIS representation
Under dual-HIS full-duplex operation, SI arises from spectral coupling between transmit and receive excitations, which is typically overlooked in conventional ISAC studies. Although the proposed Fourier-domain representation provides a promising framework for characterizing such coupling, its integration with dedicated SI suppression mechanisms remains an open problem. Moreover, the current ISAC signal model is formulated based on stochastic signal design via covariance optimization. Extending this framework toward deterministic sensing waveform design is another important direction for future investigation. At the system level, additional performance metrics, such as communication capacity, energy efficiency, and real-time computational latency, will also be explored to achieve a more comprehensive evaluation of practical performance.
-B Proof of Remark 1
We now justify the i.i.d. Gaussian model adopted in the discrete domain. Specifically, we assume the receive-side noise is a spatially white circularly symmetric complex Gaussian random field satisfying [34]
$\centering\mathbf{E}\left(v_{r}(p^{\prime})v_{r}^{*}(p^{\prime\prime})\right)=\sigma_{r}^{2}\delta\left(p^{\prime}-p^{\prime\prime}\right),~p^{\prime},p^{\prime\prime}\in D,$ (-B.1)
where $\delta(\cdot)$ is the Dirac delta function. Let $\left\{\boldsymbol{\Phi}_{n}(p^{\prime})\right\}_{n=1}^{N}$ be the truncated orthonormal Fourier basis. The $n$-th Fourier coefficient of the noise is defined as
$\centering v_{r,n}=\int_{D}v_{r}(p^{\prime})\boldsymbol{\Phi}_{n}(p^{\prime})\mathrm{d}p^{\prime},~n=1,.,N.$ (-B.2)
Collecting $v_{r,n}$ into vector $\boldsymbol{v}_{r}$, with the the orthogonality of $\boldsymbol{\Phi}_{n}(\cdot)$, we obtain
$\mathbf{E}\left(\boldsymbol{v}_{r}\boldsymbol{v}_{r}^{H}\right)=$ (-B.3)
$\left[\int_{D}\int_{D}\mathbf{E}\left(v_{r}(p^{\prime})v_{r}^{*}(p^{\prime\prime})\boldsymbol{\Phi}_{n}(p^{\prime})\boldsymbol{\Phi}_{n}^{*}(p^{\prime\prime})\right)\mathrm{d}p^{\prime}\mathrm{d}p^{\prime\prime}\right]_{n,n^{\prime}}$
$=\sigma_{r}^{2}\boldsymbol{I}_{N}.$
Therefore, $\boldsymbol{v}_{r}\sim\mathcal{CN}\left(0,\sigma_{r}^{2}\boldsymbol{I}_{N}\right)$, which validates the noise model in (23).